OpenStudy (anonymous):

Enthalpy and changes in temperature?? The specific heat of a substance is 1.72 J/g(degrees C). If given 9.52 grams of the substance and if 340 joules of heat is absorbed, what is the change in temperature of the substance? A) 4.8 x 10^-2 degrees C B) 21 degrees C C) 5600 degrees C D) 61 degrees C Can anyone help me figure out how to solve this??

OpenStudy (anonymous):

Q = m x c x change of temperature change of temperature = Q/(m x c) = 340/(1.72 x 9.52)=20.76=21